if the circumference of a circle is 16 cm .what is the area?
Answer by nyc_function(2733)

The circumference of a circle is the perimeter of the circle.
We must use the formula C = 2pir to find the radius of the circle as step one.
The circumference is given to be 16cm.
In that case, let C = 16 and solve for r. In other words, we are searching for the radius.

C = 2pir
16 = 2pir
16/2pi = r
(8/pi) = r
Now that we know the value for r (our radius), we use the formula for finding the area of the circle.
The formula for finding the area of a circle is: A = pi(r^2)
Let r = (8/pi)
A = pi(8/pi)^2
A = pi(64/(pi)^2
A = (64/pi)cm^2
Did you follow or are you still lost?